### v1

# SQ-LLaVA: Self-Questioning for Large Vision-Language Assistant

Abstract
Recent advancements in the vision-language model have shown notable generalization in vision-language tasks after visual instruction tuning. However, bridging the gap between the pre-trained vision encoder and the large language models becomes the whole network’s bottleneck. To improve cross-modality alignment, existing works usually consider more visual instruction data covering a broader range of vision tasks to fine-tune the model for question-answering, which are costly to obtain. However, the image contains rich contextual information that has been largely under-explored. This paper first attempts to harness this overlooked context within visual instruction data, training the model to self-supervised ‘learning’ how to ask high-quality questions. In this way, we introduce a novel framework named SQ-LLaVA: Self-Questioning for Large Vision-Language Assistant. SQ-LLaVA exhibits proficiency in generating flexible and meaningful image-related questions while analyzing the visual clue and prior language knowledge, signifying an advanced level of generalized visual understanding. Moreover, fine-tuning SQ-LLaVA on higher-quality instruction data shows a consistent performance improvement compared with traditional visual-instruction tuning methods. This improvement highlights the efficacy of self-questioning techniques in achieving a deeper and more nuanced comprehension of visual content across various contexts. Our implementation is available at https://github.com/heliossun/SQ-LLaVA.
1 Introduction
The recently emerging large vision-language methods, such as large language-and-vision assistant (LLaVA) and its variants Liu et al. (2023a, b); Chen et al. (2023); Bai et al. (2023); Zhang et al. (2024), fine-tune large language models (LLM) Zheng et al. (2023); Zhang et al. (2024); Bai et al. (2023) on visual instruction data to realize diverse open-world multimodal understanding, demonstrating a surprising efficacy of visual instruction tuning – the LLM learns to perform complex vision tasks simply by conditioning on a prompt containing image and text clues. Existing visual instruction data is mainly built upon conversations (e.g., ChatGPT/GPT4-V 202 (2023)), each consists of an image and multiple question-answer pairs. Building high-quality visual instruction data usually requires images and texts from different tasks to generate diverse questions, such as “Please provide the bounding box coordinate of the region this sentence describes A dead leaf on the ground.” for the object detection task. Empirically, by increasing the diversity of questions, LLaVA has achieved better performance on GQA and VizWiz tasks (26% and 45% over previous state-of-the-art methods Dai et al. (2023)). This evidence strongly suggests the advantage of training models on a broad spectrum and diverse array of tasks for enriching general vision-language understanding.
(a) Training strategy comparison.
(b) Performance comparison.
Figure 1: A high-level comparison between visual instruction tuning and visual self-questioning (ours) for vision-language assistant. SQ-LLaVA achieves state-of-the-art performance on 9 out of 10 tasks compared with other open-ended models.
LLaVA Liu et al. (2023b) model family usually consists of a pre-trained vision encoder (e.g., CLIP-ViT Radford et al. (2021)), a large generative language model (e.g., Vicuna Zheng et al. (2023), LLaMA Zhang et al. (2024), and QWen Bai et al. (2023)), and a vision-to-language projector implemented by a few linear layers. However, the modality gap between the pre-trained vision encoder and the language model restricts both sides’ generalization ability and feature representation. To overcome this challenge, various techniques have been proposed to align the vision and language domains, which could be roughly categorized into three groups: 1) build a more robust image feature extractor  Zhang et al. (2024); Zhu et al. (2024); Dai et al. (2023), 2) collect more high-quality training data Liu et al. (2023b); Chen et al. (2023); Zhu et al. (2024); Dai et al. (2023), and 3) fully fine-tune the vision and language models simultaneously during the pre-training stage Chen et al. (2023). While these methods have shown good progress in mitigating the domain gap, they inevitably bring higher computational costs and more expensive data collection and may also require sophisticated manual designs and heavy annotating efforts. Moreover, the image encompasses a wealth of information, including color, context, and the relationships between objects, etc., but most visual instruction datasets capture only a fraction. We posit that leveraging such under-explored knowledge could significantly aid in multi-modal understanding. In this study, we propose a new visual self-questioning approach by training the model to ask questions and discover vision clues without collecting extra data from other sources.
Unlike the existing visual instruction tuning methods, which treat the answer prediction as the only training target, the proposed visual self-questioning aims to mine the informative context of questions. For example, as shown in Fig. 0(a), the questions: “What is the person wearing on there feat?” and “How many bears are there?” has the answer “Socks” and “1”. Obviously, the question contains more image-related information missed by the answer, so only training the model to predict the answer under this situation has no sufficient guidance to optimize the general vision-language understanding. To this end, due to the diverse semantic information, questions within current visual instruction data can be utilized to fine-tune the instructional LLM. Therefore, we introduce a coupled visual self-questioning task besides learning the question-answering tasks.
This work proposes a novel visual self-questioning and instruction tuning model, SQ-LLaVA, utilizing the questions within the instruction data as another learnable resource to train the instructional LLMs (Fig. 0(a)). To jointly align the vision and language domains efficiently, we apply LoRAs Hu et al. (2022) to the vision encoder and the LLM and optimize them together. SQ-LLaVA proposes a prototype extractor to enhance visual representation since the learned clusters with meaningful semantic information can improve vision language alignment. Through extensive experiments (Fig. 0(b)), we show that SQ-LLaVA can further improve general vision understanding than existing visual instruction tuning methods. We summarize the contributions as follows.
  • We propose a new training technique called visual self-questioning for vision-language assistant (SQ-LLaVA) by taking advantage of informative question context in instruction data. This new SQ learning task encourages instructional LLM to capture the relationship between images and questions, leading to better vision-language alignment without collecting new data.
  • We develop a lightweight architecture of the vision-language model for SQ-LLaVA. Specifically, the prototype extractor effectively enhances the vision embedding with meaningful prototype information, and the ViT-LoRA and LLM-LoRA efficiently align vision and language domains during training.
  • Extensive experimental results show that our proposed method leads to better performance in several areas. In traditional Visual Question-answering tasks, we achieve $2.5\%$ improvement over the LLaVA-v1.5. For visual instruction tasks, our improvement is $7.2\%$. Moreover, we improve the performance by $2\%$ in cross-domain image captioning.
2 Related Work
2.1 Instruction Tuning
Instruction tuning emerged as a pivotal methodology within the realm of natural language processing (NLP), facilitating Large Language Models (LLMs) such as GPT-3 Brown et al. (2020), PaLM Chowdhery et al. (2022), and LLaMA Touvron et al. (2023a) to interpret and execute human language instructions across a spectrum of NLP tasks. This approach diverges from traditional fine-tuning mechanisms by incorporating a specialized data structure, termed instruction-following data Wei et al. (2022), which is instrumental in the fine-tuning process of LLMs. Predominantly, there are two principal methodologies concerning instruction tuning. Early scholarly contributions delineated the initial methodology, closed-domain instruction tuning Wei et al. (2022); Sanh et al. (2022); Xu et al. (2024). These studies engaged LLMs with a comprehensive assortment of publicly accessible datasets and subsequently assessed their performance across diverse NLP tasks Triantafillou et al. (2019). The empirical evidence from these inquiries consistently indicated that integrating varied NLP task instructions significantly augments the LLMs’ efficacy in navigating novel tasks. Nonetheless, LLMs calibrated with such closed-form instructions exhibited limitations in real-world user scenarios, prompting the development of an alternative approach. To address these constraints, the concept of open-domain instruction tuning Ouyang et al. (2022a); Wang et al. (2022a) was conceived. OpenAI pioneered this approach by employing human annotators to compile a corpus of real-world question-answer datasets. These datasets were the foundation for training a reward model through reinforcement learning methodologies. The trained reward model then functioned as a supervisory mechanism for further training instruction-oriented language models, such as InstructGPT Ouyang et al. (2022b) and Vicuna Zheng et al. (2023). This innovation marked a significant advancement in the field, aiming to bridge the gap between LLM performance and real-world applicability by leveraging instruction data derived from authentic user interactions.
2.2 Large Vision Language Models
As the field of LLMs and instruction tuning undergoes rapid advancements, the academic research community is progressively focusing on integrating visual information into LLM frameworks to facilitate visual instruction tuning. This emerging area of research has witnessed the proposal of various methodologies based on the foundational work of CLIP Radford et al. (2021) and diverse LLM Zheng et al. (2023); Touvron et al. (2023b); Bai et al. (2023) architectures. Notably, LLaVA Liu et al. (2023a) pioneered the integration of an LLM with a CLIP vision encoder to construct a vision-language model, demonstrating remarkable capabilities in image-text dialogue tasks through strategies of pre-training alignment and targeted instruction tuning. Subsequent investigations have sought to refine visual instruction tuning by enhancing the quality and variety of the datasets used during the pre-training and fine-tuning phases. Building upon these advancements, recent studies, including LLaVA-v1.5 Liu et al. (2023b) and ShareGPT4V Chen et al. (2023), have achieved notable success in general vision-language comprehension, showcasing their ability to undertake complex question-answering tasks. This progression underscores the importance of sophisticated data handling and model-tuning strategies in developing effective vision-language models.
3 Method
3.1 Architecture Overview
The proposed SQ-LLaVA model (see Fig. 2) consists of four main components: 1) A pre-trained vision encoder CLIP-ViT Radford et al. (2021) that extracts a sequence embedding of image tokens $Z_{v}$ for an input image $X_{v}$; 2) A prototype extractor $\phi(\cdot)$ learning visual clusters to enhance the original image tokens; 3) A trainable projection block $W(\cdot)$ with two linear layers to map the enhanced image tokens to the language domain tokens $H_{v}$, handling the dimension misalignment between the vision and language domain; and 4) Our LLM backbone $f(\cdot)$ implemented by the pre-trained Vicuna Zheng et al. (2023) to predict the next token upon the previous embedding sequence. Given the input question $X_{q}$ and answer $X_{a}$, a word embedding matrix is used to map them to contextual embeddings $H_{q}$ and $H_{a}$, and the distribution over $H_{a}^{(i+1)}$ can be obtained following the auto-regressive model as:
$p_{\theta}(H_{a}^{(i+1)}\mid H_{v},H_{q},H_{a}^{(1:i)})=\sigma(f(H_{v},H_{q},H_{a}^{(1:i)})),$ (1)
where $\theta$ represents all the trainable parameters in our model, $\sigma(\cdot)$ is a softmax function, and $f(\cdot)$ outputs the last token embedding of the whole sequence. We denote $p_{\theta}$ as the prediction probability for the anticipated answer token $H_{a}^{(i+1)}$ at the position $i+1$, conditioning on the input image token embedding $H_{v}$, the question token embedding $H_{q}$, and the previous answer token embeddings $H_{a}^{(1:i)}$. As shown in Eq. (1), the proposed SQ-LLaVA applies the language model $f(\cdot)$ to model $p_{\theta}$ given by $H_{v}$, $H_{q}$, and $H_{a}^{(1:i)}$. Existing visual instruction tuning methods Liu et al. (2023b); Chen et al. (2023) are only able to predict $H_{a}$, yet cannot fully exploit the rich semantic clues within $H_{q}$. In this study, we propose visual self-questioning instructions to guide LLMs in capturing the visual content within questions.
Figure 2: Model architecture of SQ-LLaVA. We propose prototype extractor to extract visual clustering information to enhance the visual embedding encoded by the visual encoder. SQ-LLaVA defines a new token [vusr] as specific instruction for LLM to perform visual self-questioning. Besides question answering, SQ-LLaVA treats questioning as another training objective.
3.2 Visual Self-questioning Instruction
In broad real-world scenarios, proactively asking a question requires more understanding and background knowledge than answering Tofade et al. (2013). Similarly, this work proposes visual self-questioning to encourage the LLM to discover deeper vision-language alignment and improve the overall instruction-following performance. Particularly, SQ-LLaVA treats questioning as a new learning objective beyond answering, which, to the best of our knowledge, is the first practice in the field of visual instruction tuning. While the current vision-language model can ask questions 202 (2023), such skill is still learned from question-answering through instruction tuning. However, our proposed method shows that the decoder-based LLM has the potential to learn more skills, such as how to ask questions spontaneously when a unique instruction token is given (e.g., we define [vusr] in our work). Furthermore, visual self-questioning can potentially improve general vision-language understanding. To be specific, as shown in Fig. 1, there are a certain amount of questions containing more meaningful image-related information than answers in the existing visual instruction data Liu et al. (2023b); Chen et al. (2023). Thus, we hypothesize that the vision-language understanding can be improved once the LLM learns how to predict informative questions about a given image.
Self-Questioning Prompt Design. We provide ground-truth content for the visual self-questioning, restricting SQ-LLaVA from asking image-related questions. To this end, we leverage questions as another learnable resource and follow the regular auto-regressive training objective. As shown in Fig. 3, the training data for SQ-LLaVA is designed in a format with a pre-defined template. To be specific, the system message as a fixed prompt is added at the beginning of each instruction data, indicating a general job description (e.g., gives helpful answers and asks complex questions) for the LLM. Existing visual instruction tuning methods utilize the unique tokens ([usr], [aswr]) to give the LLM a particular instruction (e.g., questioning understanding, answer prediction, etc.) and apply the delimiter token $<o^{d}>$ to mark the ending position. In this work, we propose a new special token [vusr], indicating a specific instruction – asking questions. Combining with the delimiter, we can construct instructions for self-questioning as a new training task.
Figure 3: The input sequence used to train SQ-LLaVA. Our model is trained to predict question, answer, and where to stop. We use tokens to represent learnable tokens, where $X_{q}$ is the question, $X_{a}$ is the answer, and $<o^{d}>$ is the delimiter token. In SQ-LLaVA, the System-message = "The assistant gives helpful, detailed, and polite answers to the user’s questions. Also, the assistant is a curious virtual user can ask complex questions that are relevant to the content in the image."
Each sample of current visual instruction data consists of one image $X_{v}$ and $P$ question-answer pairs $(X_{q}^{(1)},X_{a}^{(1)},\cdots,X_{q}^{(P)},X_{a}^{(P)})$. We collect one question $X_{q}^{(j)}$ and its answer $X_{a}^{(j)}$ with special tokens to construct the $j^{th}$ turn conversation as
$X_{c}^{j}=\left\{\begin{matrix}(\texttt{[usr]},X_{q}^{(j)},\texttt{[aswr]},X_{a}^{(j)})&~{}~{}j=1\mathrm{~{}or~{}}j>1,R<\delta\\ ~{}(\texttt{[{\color[rgb]{0.5625,0.66796875,0.86328125}vusr}]},X_{q}^{(j)},\texttt{[aswr]},X_{a}^{(j)})&j>1,R>\delta\end{matrix}\right.,$ (2)
where $R\in[0,1]$ is a random number and $\delta=0.5$ is a threshold that sets the proportion of self-questioning pairs in the conversations. Finally, the full-text input $X_{c}$ will be mapped to textual embedding $H_{c}$ through word embedding.
SQ-LLaVA performs zero-shot questioning without any in-context knowledge or human language instruction since the only instructional prompt is a unique token [vusr]. After visual self-questioning on instruction data with various question formats $X_{q}$, the questions sampled by SQ-LLaVA are more diversified than GPT4-V (as shown in Fig. 4), since the LLM has learned alignment between image and questions, which is different from previous works 202 (2023); Wang et al. (2022b). Specifically, previous general-purpose vision language models such as GPT4-V 202 (2023) can generate questions based on a given image, but it requires explicit language instruction such as “Ask complex questions that are relevant to the content in the image”. Accordingly, the quality of generated questions highly relies on prompt engineering. Also, self-instruct Wang et al. (2022b) utilizes in-context learning to prompt the LLM to ask specific questions.
3.3 Enhanced Visual Representation
Unlike previous visual instruction tuning methods, SQ-LLaVA jointly benefits from visual self-questioning and question-answering. For better visual self-questioning, we develop a prototype extractor that recognizes and groups similar patterns of visual information from the latent space. Our primary goal is to enhance visual representation through prototype learning.
Specifically, we utilize clustering to extract centroid representations of image tokens $Z_{v}$, where each cluster center is treated as a prototype, which, in return, will be distributed to each of the original image token embeddings. Our proposed prototype extractor $\phi(\cdot)$ is a lightweight design involving two parts: 1) cluster center optimization and 2) prototype information distribution. Following Ikotun et al. (2023); Vattani (2009), we randomly initialize $K=256$ cluster centers $C$ and deploy the iterative Expectation-Maximization (EM) clustering process to capture representative semantics in the latent space by
[…]
where $\mathcal{M}^{(t)}\in\left[0,1\right]$ denotes a soft cluster assignment map at the $t^{th}$ iteration, $\sigma(\cdot)$ is a softmax function, and $t\in\left\{1,\cdots,T\right\}$ indexes the iteration of EM step with $T=2$ in this work. Three trainable linear layers $q$, $k$, and $v$ are used in (3), where $q(\cdot)$ projects the prototype $C$ to a query vector, and $k(\cdot)$ and $v(\cdot)$ project $H_{v}$ into key and value vectors, followed by a normalization layer, respectively. The prototype extractor iteratively updates cluster map $\mathcal{M}$ and centers $C$.
After the prototype extraction, we train a linear layer $z(\cdot)$ to adaptively map the visual cluster information to the raw image embedding $Z_{v}$. For the $i^{th}$ token embedding $Z_{v}^{(i)}$, we update it as
$Z_{v}^{(i)}=Z_{v}^{(i)}+z(\frac{1}{K}\sum\nolimits_{j=1}^{K}S_{c}(C_{j},Z_{v}^{(i)})\times C_{j}),$ (4)
where $S_{c}(\cdot,\cdot)$ is a normalized cosine similarity function. The weighted sum over prototypes in Eq. (4) emerges as an indispensable step for contextual understanding from image tokens, recognizing and grouping similar patterns and semantics. It clusters image tokens as prototypes that display homogeneity in semantics, such as “grass” and “dog”. The prototypes can describe the intrinsic semantic meanings by aggregating entities that exhibit shared attributes. Finally, we map the image sequence embedding $Z_{v}$ to language domain $H_{v}$ with a two-layer linear projector $W(\cdot)$.
3.4 Model Training
3.4.1 Stage1: Pre-training for Vision-Language Alignment.
Unlike text-only LLMs, the general-purpose vision language model also fine-tunes LLM using image tokens as input (see Fig. 2). Therefore, the pre-training stage aims to optimize the LLM by explicitly executing the visual instruction. The proposed SQ-LLaVA adopts Vicuna Zheng et al. (2023) as its instruction LLM, pre-training on massive text corpora to predict the next text token given the previous context, not only containing text but also visual instructions. To achieve this, we organize the pre-training data as $D_{PT}=\{[X^{(1)}_{v},X^{(1)}_{a}],\cdots,[X^{(N)}_{v},X^{(N)}_{a}]\}$, where $N$ is the total number of training samples, and each sample has an image and its related descriptions. Each image and text input pair will be mapped to sequence embeddings ($H_{v}$ and $H_{a}$) as elaborated in Section 3.1. During pre-training, we freeze the vision encoder and LLM and mainly train the prototype extractor $\phi$ and the vision-to-language projector $W$. The pre-training goal is to maximize the probability of the predicted image description $H_{a}$ given an image $H_{v}$. When training a visual instructional LLM, we follow the negative log-likelihood objective function as
$\sum_{v,a\in D_{PT}}-\mathrm{log}p_{\theta}(H_{a}\mid H_{v})=\sum_{v,a\in D_{PT}}\sum_{i=1}^{L}-\mathrm{log}p_{\theta}(H_{a}^{(i+1)}\mid H_{v},H_{a}^{(1:i)}),$ (5)
where $L$ denotes the sequence length of answer tokens in $H_{a}$, $\theta$ is the total trainable parameter of $\phi$ and $W$, $p(H_{a}\mid H_{v})$ can be computed by Eq. (1), and $H_{a}^{(1:i)}$ represents all the answer tokens before the current prediction $H_{a}^{(i+1)}$.
3.4.2 Stage2: Fine-tuning.
Existing methods, such as LLaVA Liu et al. (2023a); Chen et al. (2023), mainly update the vision-to-language projector (usually a couple of linear layers) and the language model during fine-tuning. Nevertheless, the projector might be too weak to capture the relationship between the image and the questions. Following the previous multi-modal understanding method Radford et al. (2021), we unfreeze the vision encoder and LLM during fine-tuning for a joint optimization further to eliminate the gap between the vision and language domain.
To mitigate the heavy computational overhead, we take advantage of LoRA Hu et al. (2022) as a lightweight tuning option that can achieve similar (even better) performance to fully fine-tuning when training large models on a relatively small dataset. We add LoRA in both the vision encoder and LLM. Thus, the learnable parameters $\theta$ of the proposed SQ-LLaVA during fine-tuning represent a combination of all the parameters of LLM-LoRA, ViT-LoRA, prototype extractor $\phi$, and the vision-to-language projector $W$. Given the instruction tuning data $D_{IT}=\{[X^{(1)}_{v},X^{(1)}_{c}],\cdots,[X^{(N)}_{v},X^{(N)}_{c}]\}$, we take the conversational data $X_{c}$ and the image $X_{v}$ as input, mapping them to sequence embedding ($H_{c}$ and $H_{v}$) as elaborated in Section 3.1, and minimize the negative log-likelihood loss for the self-questioning and answering tasks as follows
[…]
where $j\in\left\{1,\cdots,P\right\}$, indicating the index of question or answer within the conversational data $X_{c}^{(*)}$. Notably, previous works Liu et al. (2023b); Chen et al. (2023); Zhu et al. (2024) only involve answering tasks, but we introduce visual self-questioning as an additional training task for visual instruction tuning. Eventually, SQ-LLaVA, as a vision-language assistant, not only executes human instructions by optimizing the objective function in Eq. (7) but can raise questions out of the given image after optimizing the Eq. (6), potentially yielding more diverse question-answer guidance and improving multi-modal understanding.
4 Experiments
4.1 Experimental Setting
4.1.1 Dataset.
Our work uses the open-source visual instruction dataset provided by LLaVA Liu et al. (2023b) and ShareGPT4V Chen et al. (2023) for training. Each dataset has large-scale image-text paired data for pre-training and instruction-following data for fine-tuning. Specifically, the instruction data proposed by LLaVA is a comprehensive mixture of COCO, GQA Hudson and Manning (2019), OCR-VQA Mishra et al. (2019), TextVQA Singh et al. (2019), VisualGenome Krishna et al. (2016), RefCOCO Kazemzadeh et al. (2014), and image from InstructBLIP Dai et al. (2023) involves multiple reasoning, spatial understanding, multi-step inference, optical character recognition, and grounding of visual concepts to language. The ShareGPT4V dataset consists of the same mixture image as LLaVA but enrolls more images from other datasets such as SAM Kirillov et al. (2023) and WebData Ordonez et al. (2011); Schuhmann et al. (2021).
[…]
Table 1: Comparison with state-of-the-art methods on ten benchmarks. After training on the same instruction data, SQ-LLaVA traind on Liu et al. (2023b) and $*{}^{*}$ trained on Chen et al. (2023) surpass their baseline model LLaVA-v1.5 and ShareGPT4V on 9 out of 10 and 6 out of 10 benchmarks in the 7B scale, and 8 out of 10 and 6 out of 10 in the 13B scale. The best results are bold and the second-best results are underlined.
[…]
7B InstructBLIP Dai et al. (2023) - 49.2 34.5 60.5 50.1 - 26.2 60.9 36.0 23.7
Qwen-VL Bai et al. (2023) 78.8 59.3 35.2 67.1 63.8 - - - 38.2 7.4
Qwen-VL-chat Bai et al. (2023) 78.2 57.5 38.9 68.2 61.5 - - - 60.6 56.7
LLaVA-v1.5 Liu et al. (2023b) 78.5 62.0 50.0 66.8 58.2 85.9 30.5 63.4 64.3 58.3
ShareGPT4V Chen et al. (2023) 80.6 63.3 57.2 68.4 60.4 86.8 37.6 72.6 68.8 62.2
SQ-LLaVA 79.2 62.8 54.0 68.9 58.6 87.7 32.5 66.3 66.2 58.1
$*{}^{*}$ 80.3 63.7 55.3 70.5 60.5 87.2 37.6 74.3 66.6 60.0
13B InstructBLIP Dai et al. (2023) - 49.5 33.4 63.1 50.7 78.9 25.6 58.2 - -
LLaVA-v1.5 Liu et al. (2023b) 80.0 63.3 53.6 71.6 58.2 85.9 35.4 70.7 67.7 63.6
ShareGPT4V Chen et al. (2023) 81.0 64.8 55.6 71.2 62.2 - 43.1 79.9 68.5 63.7
SQ-LLaVA 80.1 63.6 54.6 69.8 60.2 87.7 35.5 74.6 68.7 62.0
$*{}^{*}$ 81.3 65.0 58.2 71.5 61.9 87.4 39.7 80.7 68.5 62.5
Figure 4: Visual self-questioning of SQ-LLaVA-7B. Comparing to the question data provided by GPT4-V (data collected by LLaVA-v1.5 Liu et al. (2023b)), SQ-LLaVA can generate questions with higher diversity i.e. multiple choice and tricky questions.
We also evaluate our model for visual information discovery through captioning. To be specific, we employ testing images from four datasets, i.e., COCO Chen et al. (2015), Flickr30K Young et al. (2014), Conceptual Sharma et al. (2018), and Nocaps Agrawal et al. (2019). For Flickr30K, we adopt the standard split and focus on its test set, which consists of $1000$ samples. We use Nocaps-out-domain split, which exclusively contains images depicting objects absent in COCO. Following previous research Dessì et al. (2023), we use the Nocaps validation set for testing. Following Dessì et al. (2023), we evaluate the proposed method with regular image captioning metrics, e.g., BLEU Papineni et al. (2002) and CIDEr Vedantam et al. (2015). For the open-world methods LLaVA-v1.5 Liu et al. (2023b), ShareGPT4V Chen et al. (2023) and the proposed SQ-LLaVA, we utilize greedy search for caption generation with a prompt of “Provide a brief description of the given image, your answer should be in one sentence.” The generated caption is used to evaluate the performance of image captioning.
4.1.2 Compared Methods.
In this work, we choose baseline models with similar architecture. Precisely, each model consists of a vision encoder and an LLM with different model size and pre-trained weights. For both image description and question answering tasks, we adopt InstructBLIP Dai et al. (2023), Qwen-VL Bai et al. (2023), LLaVA-v1.5 Liu et al. (2023b), and ShareGPT4V Chen et al. (2023) as our baseline models, where all methods were trained on visual instruction data. Besides, we also incorporate ClipCap Mokady (2021) and DiscriTune Dessì et al. (2023), which are trained on regular image-text pair data.
4.1.3 Implementation.
We adopt Vicuna Zheng et al. (2023) as the pre-trained language generative model and CLIP-ViT Radford et al. (2021) as the vision encoder. We pre-train the prototype extractor and the vision-to-language projector using AdamW Loshchilov and Hutter (2019) optimizer with a learning rate of $2\times 10^{-3}$ and a constant scheduler for one epoch. Following previous work Liu et al. (2023b), we keep the global batch size as $256$ for pre-training and $128$ for fine-tuning. During fine-tuning, we insert LoRAHu et al. (2022) with $rank=128$ and $\alpha=256$ into the language model (LLM-LoRA) and LoRA with $rank=32$ and $\alpha=64$ into the vision encoder (ViT-LoRA). We optimize LoRA modules, the prototype extractor, and vision-to-language projector for one epoch. We set the learning rate to $2\times 10^{-4}$ for LoRA, and $2\times 10^{-5}$ for the other layers.
Table 2: Image captioning comparison on four datasets. All methods are trained with images from COCO. SQ-LLaVA is trained on data collected by Liu et al. (2023b). $*{}^{*}$ is trained on data collected by Chen et al. (2023). We adopt B@4 and CIDEr for evaluation.
[…]
B@4 CIDEr B@4 CIDEr B@4 CIDEr B@4 CIDEr
ClipCap Mokady (2021) 2B 8.50 37.03 17.21 41.65 20.32 51.74 1.47 23.74
DiscriTune Dessì et al. (2023) 2B 13.99 53.20 18.48 44.78 24.10 57.06 1.71 28.01
LLaVA-v1.5Liu et al. (2023b) 7B 30.68 112.94 28.67 81.27 35.78 103.56 2.79 39.20
ShareGPT4V Chen et al. (2023) 7B 30.54 110.71 31.00 86.17 37.19 107.45 2.78 37.86
SQ-LLaVA 7B 29.85 110.77 29.88 83.51 36.21 105.42 2.90 39.49
$*{}^{*}$ 7B 30.13 109.71 31.49 83.14 37.20 107.42 2.91 41.24
Figure 5: A qualitative evaluation of detailed and brief image descriptions from three models. We highlight the words and sentences that represent how each model describes the main object in the image.
4.2 Zero-shot Multilingual Capability
We evaluate SQ-LLaVA on ten benchmarks, covering a range of academic Visual Question Answering (VQA) tasks and recent instruction tuning tasks designed for large vision language models. The academic VQA includes VQA-v2 Goyal et al. (2017) and VizWiz Gurari et al. (2018). GQA Hudson and Manning (2019) is a fine-grained real-world visual reasoning and question-answering benchmark. ScienceQA Lu et al. (2022) is a benchmark with rich subjects (natural science, language science, and social science). TextVQA Singh et al. (2019) requires the model to recognize the texts in the image. LLaVA (in the wild) and MM-Vet Yu et al. (2023) use GPT4 to assess the capability of the models for testing. With manually designed questions, MM-Bench and MMBench-CN Liu et al. (2023c) evaluate the model’s vision-related reasoning and perception for English and Chinese, respectively. POPE Li et al. (2023) is a benchmark for evaluating object hallucination Rohrbach et al. (2018).
In Table 1, we quantitatively compare between SQ-LLaVA and existing models. SQ-LLaVA-7B and SQ-LLaVA-13B trained on two instruction datasets Liu et al. (2023b); Chen et al. (2023) outperform previous methods in six out of ten visual instruction tuning tasks. To be specific, SQ-LLaVA-7B achieves $17.2\%$ improvement over LLaVA-v1.5-7B on the LLaVA (in the wild) benchmark, indicating the superior capabilities of our model in tasks such as detailed description and complex reasoning. Also, SQ-LLavA-7B improves over previous methods on ScienceQA, indicating that our model excels in understanding and reasoning over scientific content and can effectively handle multi-modal information. The improvement in ScienceQA suggests strong capabilities in multi-hop reasoning, comprehension of complex scientific concepts, and the ability to utilize context and explanations to derive correct answers. SQ-LLaVA-7B has a steady improvement over LLaVA-v1.5-7B and ShareGPT4V-7B on the POPE benchmark, and the $2\%$ and $1\%$ improvement indicates that our proposed method has better reliability and trustworthiness since POPE is a task designed to evaluate the phenomenon of object hallucination Li et al. (2023); Rohrbach et al. (2018). In the bottom section of Table 1, the proposed SQ-LLaVA-13B surpasses previous works in six out of ten benchmarks, indicating the scalability of our method on larger LLM. Notably, all the improvements are achieved with significantly fewer trainable parameters compared to other methods Chen et al. (2023); Liu et al. (2023b); Bai et al. (2023). To further evaluate SQ-LLaVA, we present experiments on one recent benchmark in Appendix A.2.
4.3 Visual Information Discovery
In this experiment, we showcase the diversity and reliability of the proposed SQ-LLaVA through various qualitative applications, including detailed image description, visual information summary, and visual self-questioning. We also present quantitative results on the task of image captioning.
Abilities of SQ-LLaVA Through Qualitative Samples. SQ-LLaVA demonstrates many advanced abilities compared to traditional vision language models. For example, SQ-LLaVA can effectively eliminate object hallucination Li et al. (2023); Rohrbach et al. (2018), delivering the predictions with better trustworthiness. Besides, SQ-LLaVA can generate diverse and meaningful questions about a given image without human textual instructions.
In Fig. 4, we show visual self-questioning results of SQ-LLaVA. We find SQ-LLaVA can sample questions with high diversity and quality. For example, the generated questions include multiple-choice, reasoning, binary choice questions, and the content within each question, all of which are either highly correlated or tricky to the given image, such as “What is the hair color of the man that is looking up at the ocean?” from the third image and “Are there cows in the picture?” from the second image.
The detailed image description example in Fig. 5 demonstrates that SQ-LLaVA can generate visual information with concrete concepts within the image, such as “Hyundai”, “small white dog”, “Nikky Stephen”. In contrast, LLaVA-v1.5 can only describe the image from a general concept, such as “car”, “dog”, or “everyday moment”. Even though ShareGPT4V can generate detailed descriptions, it still suffers from the issue of object hallucination, such as “perhaps waiting” and “behind the duck sign”. By observation, SQ-LLaVA can describe the image with less unintended text, yielding higher reliability. On the other hand, all methods can generate similar brief image descriptions containing only the main objects within the image, indicating the promising summary skills of each model. See Appendix B for more qualitative evaluation on image description and visual self-questioning.
[…]
Table 3: Ablation study of training strategy on visual instruction tasks. All models are in 7B scale with three components of ViT-LoRA (V-LoRA), self-questioning (SQ), and prototype extractor (Proto). We provide experiments on two instruction datasets Liu et al. (2023b); Chen et al. (2023) with different pre-training (PT) and instruction tuning (IT) data scales.
[…]
558K 665K 49.4 68.4 58.2 86.5 67.1 65.9
52.4 67.9 58.6 87.7 65.6 66.4
52.6 68.4 57.8 88.2 67.3 66.9
53.4 69.3 58.1 87.9 67.9 67.3
54.0 68.9 58.6 87.7 68.1 67.5
1200K 700K 51.5 68.9 58.9 86.8 72.1 67.6
54.0 68.9 60.2 87.2 71.6 68.4
55.4 69.2 59.5 86.8 77.3 69.6
54.2 70.3 60.5 87.5 72.7 69.0
55.3 70.5 60.5 87.2 74.3 69.6
4.4 Ablation Study
In Table 3, we conduct experiments with different architecture designs and training strategies on five question-answering benchmarks. For a fair comparison, we train the baseline models on our local machine with the same training recipe of LLaVA-LoRA Liu et al. (2023b). Specifically, we present our full model and three ablated models by removing one component each time, where all models are in 7B scale. We adopt the dataset Liu et al. (2023b) with $558$k for pre-training (PT) and $665$k for fine-tuning (FT). As compared, self-questioning (SQ) brings a consistent performance boost on all the five benchmarks, indicating the effectiveness of visual self-questioning on improving visual language understanding. Besides, we introduce the prototype extractor (Proto) to enhance visual representation, achieving $0.9\%$ improvement in average accuracy among five benchmark. With all three components incorporated, we observe a $2.4\%$ improvement in average accuracy.
As shown by the bottom block of Table 3, we conduct experiments with the same ablation settings but with a larger scale of the visual instruction data Chen et al. (2023) (i.e., for both PT and IT). Overall, SQ-LLaVA achieves $2.4\%$ improvement over the baseline model after training on the smaller dataset and achieves $3.0\%$ improvement after training on the larger dataset. We also provide ablation study of SQ-LLaVA in 13B scale (see Appendix A.1 for detail).
5 Conclusion
This work introduced SQ-LLaVA, a method to improve general-purpose vision language understanding and image-oriented question answering through the proposed visual self-questioning technique. Our experiments reveal that SQ-LLaVA achieved superior performance with fewer training parameters. We performed a comprehensive study on visual instruction tuning tasks and found that SQ-LLaVA can generalize well to a wide range of unseen tasks and outperform state-of-the-art methods. Notably, qualitative assessments demonstrated that SQ-LLaVA allows enhanced visual representation and domain alignment capability, effectively alleviating the issue of object hallucination and improving the semantic interpretation of images. Our findings underscored the potential of visual self-questioning as a potent training strategy within the field of visual instruction tuning, paving the way for more efficient and effective model training approaches. SQ-LLaVA is a foundational step toward enhancing general vision language understanding by framing question-asking as an intrinsic goal of the question-answering process. We hope the proposed SQ-LLaVA can encourage future endeavors in exploiting questions in visual instruction tuning by following the proposed self-questioning technique.
Appendix A Zero-shot Multilingual Capability
A.1 Ablation Study
In Table 4, we conduct experiments with different architecture designs and training strategies on five benchmarks for SQ-LLaVA-13B. For a fair comparison, we train the baseline models on our local machine with the same training recipe of LLaVA-LoRA Liu et al. (2023b). Specifically, we present the baseline model, our full model, and three ablated models by removing one component each time. We adopt the dataset Liu et al. (2023b) with $558$k for pre-training (PT) and $665$k for fine-tuning (FT). As compared, self-questioning (SQ) achieves $0.6\%$ improvement in average accuracy, indicating the effectiveness of visual self-questioning in improving visual language understanding. Besides, we introduce the prototype extractor (Proto) to enhance visual representation, achieving $0.6\%$ improvement in average accuracy. With all three components incorporated, we observe a $2.5\%$ improvement in average accuracy among five benchmarks.
As shown by the bottom block of Table 4, we conduct experiments with the same ablation settings but with a larger scale of the visual instruction data Chen et al. (2023) (i.e., for both PT and IT). Overall, SQ-LLaVA achieves $2.5\%$ improvement over the baseline model after training on the smaller dataset and achieves $3.2\%$ improvement after training on the larger dataset.
In Table 4, we observe that the proposed SQ-LLaVA achieves significant improvement over the baseline models on the LLaVA (in-the-wild) benchmark by $7.3\%$ and $5.6\%$. This indicates SQ-LLaVA’s capability in more challenging tasks, generalizability to novel domains, and robustness to different prompts.
Table 4: Ablation study of training strategy on visual instruction tasks. All models are in 13B scale with three components of ViT-LoRA (V-LoRA), self-questioning (SQ), and prototype extractor (Proto). We provide experiments on two instruction datasets Liu et al. (2023b); Chen et al. (2023) with different pre-training (PT) and instruction tuning (IT) data scales.
[…]
558K 665K 53.9 68.2 60.2 86.7 69.5 67.7
54.5 70.4 60.6 87.7 71.6 69.0
54.8 69.4 60.7 87.5 72.8 69.0
54.1 68.6 60.1 88.0 71.0 68.4
54.6 69.8 60.2 87.7 74.6 69.4
1200K 700K 53.6 70.1 62.0 87.0 76.4 69.8
56.8 71.0 61.9 87.4 79.6 71.3
54.3 71.1 61.7 86.9 77.0 70.2
56.2 70.2 61.4 86.9 83.6 71.7
58.2 71.7 62.0 87.4 80.7 72.0
A.2 Evaluation on An Additional Benchmark
We evaluate SQ-LLaVA on one recent benchmark MME Fu et al. (2023). MME is designed to assess the performance of vision-language models across a wide range of tasks that require both perception and cognition abilities. In Table 5, we provide a comparison between SQ-LLaVA with other methods. As can be seen, SQ-LLaVA achieves the highest cognition score among all baseline models, indicating a strong performance when handling the commonsense reasoning, numerical calculation, text translation, and code reasoning tasks.
Table 5: Comparison with SOTA methods on recent evaluation benchmarks. The best results are bold and the second-best results are underlined.
[…]
InstructBLIP Dai et al. (2023) 1212.8 291.8
LLaVA-v1.5 Liu et al. (2023b) 1531.3 295.4
ShareGPT4V Chen et al. (2023) 1618.7 303.2
SQ-LLaVA 1568.2 326.8
Appendix B Qualitative Analysis of SQ-LLaVA
B.1 Reliable Detailed Image Description
The detailed image description examples demonstrate that SQ-LLaVA can capture low-level visual information within the image, such as “KL8-Q17” in Fig. 6 and “number 186 206-7” in Fig. 7. Also, SQ-LLaVA demonstrates a better reliability. In Fig. 7, the image shows a complex environment with various objects, which makes it easy for the models to generate wrong descriptions such as “total five people” (LLaVA-v1.5) and “faces blurred” (ShareGPT4V). In contrast, SQ-LLaVA describes uncertainty information with a prefix hedge word such as “likely”, “possibly”, and “perhaps”, achieving higher reliability in a real-world application.
B.2 Visual Self-questioning
In Fig. 10, Fig. 11, and Fig. 12, we show visual self-questioning and answering of SQ-LLaVA. As can be seen, SQ-LLaVA asks meaningful questions based on comprehensive visual understanding. Specifically, the self-asked questions always involve one or more main objects within the image and a relatively divergent problem, where the main objects rely on visual perception, and the meaningful problem requires vision-language alignment. Unlike question-answering with direct instruction, self-questioning encourages the model to discover all possible combinations between visual and semantic information.
Figure 6: A qualitative evaluation of detailed image descriptions from four models. We highlight the words and sentences that represent how each model describes the main object in the image. We use underlined word to indicate how SQ-LLaVA describe uncertain information.
Figure 7: A qualitative evaluation of detailed image descriptions from four models. We highlight the words and sentences that represent how each model describes the main object in the image. We use underlined word to indicate how SQ-LLaVA describe uncertain information.
Figure 8: A qualitative evaluation of detailed image descriptions from four models. We highlight the words and sentences that represent how each model describes the main object in the image. We use underlined word to indicate how SQ-LLaVA describe uncertain information.
Figure 9: A qualitative evaluation of detailed image descriptions from four models. We highlight the words and sentences that represent how each model describes the main object in the image. We use underlined word to indicate how SQ-LLaVA describe uncertain information.
Figure 10: Visual self-questioning and answering of SQ-LLaVA.
Figure 11: Visual self-questioning and answering of SQ-LLaVA.
Figure 12: Visual self-questioning and answering of SQ-LLaVA.
Figure 13: Visual self-questioning and answering of SQ-LLaVA.
Figure 14: Visual self-questioning and answering of SQ-LLaVA.